# Wukong’s 72 Transformations: High-fidelity Textured 3D Morphing via Flow Models

### v3

@@ Appendix B Ablation study
To ensure a fair, apples-to-apples comparison with 3DRM, we implemented our full morphing method on top of the GaussianAnything (yushi2025gaussiananything) framework—the same 3D generator used in 3DRM (yang2025textured), results are shown in Table LABEL:tab:ablation_gaussian. We denote this variant as “Ours*” in the table below. Across all evaluation metrics, including FID, PPL, V-CLIP, and GPT-based perceptual scores (STP-GPT, SEP-GPT), our method consistently outperforms 3DRM, even when both share the exact same backbone. This clearly demonstrates that the improvement is not solely due to the use of a stronger generator like Trellis, but rather stems from our core morphing algorithm. Note that the GPT-based results may differ from those presented in main paper, as they are computed using only the four methods reported here.
@@
Besides, We conduct quantitative evaluations with different threshold $\tau$ values and present the results below as shown in Table LABEL:tab:ablation_threshold. We observe that the performance is robust across a reasonable range of thresholds (0.2-0.8). We set a default threshold 0.3 in our evaluation.
@@
To evaluate the method’s generalization to real 3D data, we conducted experiments using the Headspace dataset (dai2020statistical), which contains high-quality 3D face scans along with corresponding rendered RGB images. In our pipeline, we used these rendered images as inputs and passed them through the DINOv2 (oquab2023dinov2) encoder to extract texture and semantic features for morphing. The outputs were generated by our standard pipeline without any architectural changes or fine-tuning. Despite relying on pretrained components, our method shows strong generalization to real-world 3D scans. Our method outperformed both MorphFlow (tsai2022multiview) and 3DRM (Our own implementation) (yang2025textured) on the same evaluation protocol. Quantitative results are shown below in Table LABEL:tab:ablation_head:
@@
There are theoretical guarantees for flow models like the rectified flow model in maintaining the convexity of data during the generation process. This linearity ensures that any intermediate sample lies within the convex hull of the endpoints, thereby preserving the convexity of the data. In practice, the trajectory is hard to remain straight. There is analysis (liu2022flowstraightfastlearning) on the straightness error on the trajectory, which states that even an imperfect trajectory is close enough to straight lines and ensure convexity of data to some extent. Theorems in this analysis further emphasize the uniqueness and optimality of the solution of rectified flow in matching distributions under convex cost functions. For diffusion models, the backward process generates data from the prior but does not theoretically guarantee convexity preservation. These models focus on matching data distributions, not preserving geometric properties like convexity. There is no theatrical guarantee on the data convexity in the backward process. The noise term in the reverse-time SDE can easily violate the convexity of original data. Also, under the same condition as the rectified flow model, the path of diffusion models is not assured to be optimal. There exists certain crossing flows in the matching of two distributions, leading to features that are out-of-distribution in practice.
@@ D.2 Comparison with shape morphing methods
Although previous 3D shape morphing methods do not consider texture transformation, we provide a comparison focused solely on shape deformation. As shown in Table LABEL:tab:appendix_classic, we compare our method with several state-of-the-art shape morphing approaches, including MapTree (ren2020maptree), BIM (kim2011blended), SmoothShells (eisenberger2020smooth), NeuroMorph (eisenberger2021neuromorph), and SRIF (sun2024srif). Following the evaluation protocol from (sun2024srif), we use the SHREC07 (temerinac2007shrec) dataset and report performance using Dirichlet energy (ezuz2019reversible) and Coverage (huang2017adjoint) metrics. Our method achieves superior results across both metrics, demonstrating more efficient and accurate shape interpolation.
For qualitative comparison, we show results against SRIF in Fig. 10, where rows 1, 3 and 5 show SRIF’s outputs and rows 2, 4 and 6 show ours. Our morphing process is smoother and preserves finer details in intermediate shapes—for example, the gecko’s toes in row 2 and the head structure in row 4. Additional comparison with NeuralMorph is presented in Fig. 11, where our results are again significantly more detailed and coherent.
@@ Appendix E More results
Fig. 12 and Fig. 13 illustrate the textured 3D morphing process generated from “Wukong” to a variety of objects. The texture can be flexibly inherited from either the source or the target image, depending on the user preference. Fig. 14 and Fig. 15 further demonstrate morphing between additional object pairs, showcasing the versatility of our method. Notably, our method is capable of performing textured 3D morphing not only between geometrically complex objects, but also across different semantic categories, highlighting its superior robustness and generalizability.
@@ Appendix G Limitation
While our method achieves state-of-the-art performance in textured 3D morphing, several limitations remain. First, like existing morphing methods, our approach still encounters difficulties on cases involving extreme topological changes, such as splitting or merging parts. These scenarios remain a general challenge in the field and are not yet fully addressed by existing methods. Second, since our method operates without explicit 3D supervision or correspondence annotations, its results may be sensitive to ambiguities in the input prompts or inconsistencies in multi-view generation, especially when the input lacks structural clarity.

### v4

@@ Appendix B Ablation study
To ensure a fair, apples-to-apples comparison with 3DRM, we implemented our full morphing method on top of the GaussianAnything (yushi2025gaussiananything) framework—the same 3D generator used in 3DRM (yang2025textured), results are shown in Tab LABEL:tab:ablation_gaussian. We denote this variant as “Ours*” in the table below. Across all evaluation metrics, including FID, PPL, V-CLIP, and GPT-based perceptual scores (STP-GPT, SEP-GPT), our method consistently outperforms 3DRM, even when both share the exact same backbone. This clearly demonstrates that the improvement is not solely due to the use of a stronger generator like Trellis, but rather stems from our core morphing algorithm. Note that the GPT-based results may differ from those presented in the main paper, as they are computed using only the four methods reported here.
@@
Besides, we conduct quantitative evaluations with different threshold $\tau$ values and present the results below as shown in Tab LABEL:tab:ablation_threshold. We observe that the performance is robust across a reasonable range of thresholds (0.2-0.8). We set a default threshold 0.3 in our evaluation.
@@
To evaluate the method’s generalization to real 3D data, we conducted experiments using the Headspace dataset (dai2020statistical), which contains high-quality 3D face scans along with corresponding rendered RGB images. In our pipeline, we used these rendered images as inputs and passed them through the DINOv2 (oquab2023dinov2) encoder to extract texture and semantic features for morphing. The outputs were generated by our standard pipeline without any architectural changes or fine-tuning. Despite relying on pretrained components, our method shows strong generalization to real-world 3D scans. Our method outperformed both MorphFlow (tsai2022multiview) and 3DRM (Our own implementation) (yang2025textured) on the same evaluation protocol. Quantitative results are shown below in Tab LABEL:tab:ablation_head:
@@
There are theoretical guarantees for flow models like the rectified flow model in maintaining the convexity of data during the generation process. This linearity ensures that any intermediate sample lies within the convex hull of the endpoints, thereby preserving the convexity of the data. In practice, the trajectory is hard to remain straight. There is analysis (liu2022flowstraightfastlearning) on the straightness error on the trajectory, which states that even an imperfect trajectory is close enough to straight lines and ensures the convexity of data to some extent. Theorems in this analysis further emphasize the uniqueness and optimality of the solution of rectified flow in matching distributions under convex cost functions. For diffusion models, the backward process generates data from the prior but does not theoretically guarantee convexity preservation. These models focus on matching data distributions, not preserving geometric properties like convexity. There is no theoretical guarantee on the data convexity in the backward process. The noise term in the reverse-time SDE can easily violate the convexity of the original data. Also, under the same condition as the rectified flow model, the path of diffusion models is not assured to be optimal. There exist certain crossing flows in the matching of two distributions, leading to features that are out-of-distribution in practice.
@@ D.2 Comparison with shape morphing methods
Although previous 3D shape morphing methods do not consider texture transformation, we provide a comparison focused solely on shape deformation. As shown in Tab LABEL:tab:appendix_classic, we compare our method with several state-of-the-art shape morphing approaches, including MapTree (ren2020maptree), BIM (kim2011blended), SmoothShells (eisenberger2020smooth), NeuroMorph (eisenberger2021neuromorph), and SRIF (sun2024srif). Following the evaluation protocol from (sun2024srif), we use the SHREC07 (temerinac2007shrec) dataset and report performance using Dirichlet energy (ezuz2019reversible) and Coverage (huang2017adjoint) metrics. Our method achieves superior results across both metrics, demonstrating more efficient and accurate shape interpolation.
For qualitative comparison, we show results against SRIF in Fig. 10, where rows 1, 3, and 5 show SRIF’s outputs and rows 2, 4, and 6 show ours. Our morphing process is smoother and preserves finer details in intermediate shapes—for example, the gecko’s toes in row 2 and the head structure in row 4. Additional comparison with NeuralMorph is presented in Fig. 11, where our results are again significantly more detailed and coherent.
@@ Appendix E More results
Fig. 12 and Fig. 13 illustrate the textured 3D morphing process generated from “Wukong” to a variety of objects. The texture can be flexibly inherited from either the source or the target image, depending on the user’s preference. Fig. 14 and Fig. 15 further demonstrate morphing between additional object pairs, showcasing the versatility of our method. Notably, our method is capable of performing textured 3D morphing not only between geometrically complex objects, but also across different semantic categories, highlighting its superior robustness and generalizability.
@@ Appendix G Limitation
While our method achieves state-of-the-art performance in textured 3D morphing, several limitations remain. First, like existing morphing methods, our approach still encounters difficulties in cases involving extreme topological changes, such as splitting or merging parts. These scenarios remain a general challenge in the field and are not yet fully addressed by existing methods. Second, since our method operates without explicit 3D supervision or correspondence annotations, its results may be sensitive to ambiguities in the input prompts or inconsistencies in multi-view generation, especially when the input lacks structural clarity.